# Event Topology Classifiers at the Large Hadron Collider

### v1

@@ Abstract
Event classifiers are the most fundamental observables to probe the system size and event topology of high-energy collisions. Significant progress has been made to establish suitable event classifiers to probe different physics processes occurring in elementary $e^{+}e^{-}$ to heavy-ion collisions in a range of center of mass energies. One of the major motivations to revisit event classifiers at the Large Hadron Collider (LHC) originates from the recent measurements of high multiplicity proton-proton collisions, which have revealed that these small collision systems exhibit features similar to the formation of quark-gluon plasma (QGP), traditionally believed to be only achievable in heavy nucleus-nucleus collisions at ultra-relativistic energies. To pinpoint the origin of these QGP-like phenomena with substantially reduced auto-correlation and selection biases, and to bring all collision systems on equal footing, along with charged-particle multiplicity, lately several event topology classifiers such as transverse sphericity, transverse spherocity, relative transverse activity classifier, and charged-particle flattenicity have been used extensively in experiments as well as in the phenomenological front. In this review article, we extensively discuss the purpose, coverage, and usage guidelines of these event shape classifiers, which can be crucial given the upcoming measurements in a precision era during Runs 3 and 4 of the LHC. In most cases, the event shape observables are found to be better probes in understanding the heavy-ion-like behavior seen at the LHC, while making a multi-differential study of multihadron production dynamics in hadronic and nuclear collisions.
I Introduction
@@
To understand the infancy of our Universe and its evolution, the little bangs of heavy-ions are made in collider experiments like the Large Hadron Collider (LHC) at CERN, Switzerland, and the Relativistic Heavy-Ion Collider (RHIC) at BNL, USA, which reproduces the conditions that happened during the fraction of a second after the Big-Bang. The collision of two nuclei produces a hot and dense system of deconfined quarks and gluons for a short time (order of a few fm/c or 10-23 sec). This deconfined and thermalized state of quarks and gluons is often referred to as quark-gluon plasma (QGP). Due to a very short timescale, it is impossible to probe QGP directly. However, several indirect signatures are proposed, such as the presence of strangeness enhancement, quarkonia suppression, jet quenching, collective flow, direct photon production, and direct dilepton production, which validate the presence of QGP. In proton-proton ($pp$) collisions, where the system size is significantly smaller than Pb-Pb, the main hard parton-parton scatterings produce high transverse momentum ($p_{\rm T}$) particles which can be described by the perturbative quantum chromodynamics (pQCD). In the hard regime of QCD, jets are produced from parton scatterings at large $p_{\rm T}$. As in this regime, the coupling constant ($\alpha_{s}$) is significantly small, precise pQCD calculations can be performed. The semi-hard and soft parton scatterings, which are not a part of the main hard scatterings, form the underlying event (UE) [1, 2]. As the soft-QCD calculations are cumbersome due to a larger value of $\alpha_{s}$, the description of the UE is modeled using different QCD-inspired phenomenological approaches [3]. At the LHC energies, due to high parton densities and the composite nature of hadrons, more than one binary interaction among partons is possible in a single $pp$ collision, which is often referred to as multi-parton interactions (MPI) [4]. The MPI phenomenon has been widely supported by experimental data [5, 6] and is one of the key ingredients in the Monte Carlo event generators. In such MPI-based models, the final state in $pp$ collisions is sensitive to the modeling of MPI and non-perturbative effects in the final state, such as color reconnection (CR) and color Ropes, which are modeled in PYTHIA 8 [7] (See Figs. 1 and 2).
@@
Figure 2: Pictorial representation of Color Reconnection (CR) in the string fragmentation model. (a) two independent hard scatterings, (b) color reconnected string
@@
Recent measurements of high multiplicity $pp$ collisions at the LHC energies have revealed that these systems exhibit features similar to quark-gluon plasma, such as the presence of radial and elliptic flow [8, 9, 10, 11, 12, 13, 14, 15, 16] and enhanced production of strange hadrons with respect to that of charged pions [17, 18, 19], traditionally believed to be only achievable in ultra-relativistic nucleus-nucleus collisions. These measurements may point towards a common underlying physics mechanism across collision systems. However, no jet quenching signatures have been reported in small collision systems [20, 21]. The event generators currently available with relevant physics models fail to simultaneously reproduce all the observed phenomena in small collisions. This makes the origin of QGP-like phenomena observed in small systems like $pp$ and p-Pb unclear. Out of several theoretical models, the explanations from the MPI-based models with final state non-pQCD phenomena like CR [7], rope hadronisation [22, 23], and string shoving [24], inherited in PYTHIA8, are able to explain many of the QGP-like signatures in $pp$ collisions. MPI is a phenomenological phenomenon and can not be measured in experiments. Thus, many experimental measurements are performed as a function of charged particle multiplicity in the final state, which has a strong correlation with MPI in MC models. On the other hand, the measurements with charged particle multiplicity show a stronger than linear increase of high-$p_{\rm T}$ particle yields in high-multiplicity (HM) $pp$ collisions relative to that of minimum bias (MB) events. This signifies a selection bias towards local fluctuations of choosing only hard $pp$ collisions. Such biases can be removed when the event selection is performed with multiplicity measured in a different pseudorapidity interval with respect to the observable of interest. However, it is found that such measurements are still biased by the hard processes contributing to the high-$p_{\rm T}$ particles [26]. Thus, in small collision systems, such selection biases in the measurements hinder the search for the origin of QGP-like behavior.
@@
To pindown the origin of QGP-like phenomena with substantially diminished selection bias and to bring all the collision systems in an equal footing, along with charged-particle multiplicity ($N_{\rm{ch}}$), lately event shape observables such as transverse spherocity ($S_{0}$) and transverse sphericity ($S_{\rm T}$) are inherited from the usage in $e^{+}e^{-}$ collisions since late 1970s [27, 28, 29, 30]. Additionally, there are other proposed event shape observables which are used in heavy-ion collisions and jet studies such as third-jet resolution parameter, $y23$ [31, 32, 33]; aplanarity, $A$; transverse thrust, $\tau_{T}$; it’s minor component, $T_{\perp,\rm min}$  [34, 35, 36]; $\mathcal{F}$-parameter [37], reduced flow vector, $q_{2}$ [38]; and event isotropy [39]. Moreover, several new event shape observables have been recently constructed, for example, relative transverse activity classifiers such as $R_{\rm T}$, $R_{\rm T}^{\rm min}$, and $R_{\rm T}^{\rm max}$ developed by P. Skands et. al. and charged-particle flattenicity ($\rho_{\rm ch}$) developed by A. Ortiz et. al. to reduce the sensitivity to hard processes compared to the classifiers measured based only on charged-particle multiplicity [40, 41, 42, 43]. In both experiments as well as in the phenomenological fronts, extensive explorations using the above-mentioned event-shape observables have been performed [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 37, 38, 39, 40, 41, 42, 43, 44, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 45, 36]. Also, we have explored the particle production dynamics as a function of transverse spherocity for light, strange [56, 57, 58, 59, 60] and heavy flavor [61, 62] sector in proton-proton collisions. The possibility of thermalization in small collision systems has been explored using spherocity as an event-shape observable in Ref. [63, 64]. The global observables, identified particle production, and anisotropic flow coefficients are studied as a function of transverse spherocity in Pb–Pb collisions in Refs. [65, 66, 67, 68, 69, 70, 71].
@@
In this review article, we extensively discuss the purpose, coverage, and usage guidelines of these event shape classifiers, which can be crucial in view of the upcoming measurements in the precision era at the LHC. Since 1970, there have been several event classifiers proposed and studied. We will mostly highlight the event classifiers which are popular in the early 21st century since the commencement of the LHC in view of the studies of small collision systems at the LHC. We will briefly discuss the most fundamental event classifier for the high energy collisions, i.e, the charged-particle multiplicity measured both in mid and forward rapidities. Then the discussion will be followed by transverse sphero(i)city, which uses the information of charged-particle transverse momenta and the azimuthal angle for each particle. Then, we will discuss the event shape classifiers such as $R_{\rm T}$, $R_{\rm T}^{\rm min}$, and $R_{\rm T}^{\rm max}$, which take the advantage of the underlying event and uses the information of charged-particle multiplicity in the transverse region and azimuthal angle of the particles. Last but not least, the study will be focused on a relatively new event shape classifier, charged-particle flattenicity, which takes the information of transverse momentum, pseudorapidity, and azimuthal angle of final state particles. Along with their novelty, the above-discussed event classifiers show significant correlation with $N_{\rm mpi}$, which will be extensively discussed in the next section. In most cases, the event shape observables are found to be better probes in understanding the heavy-ion-like behavior seen at the LHC. However, the coverage of event shape classifiers usually overlaps with each other. Thus, one needs to have a better understanding of these classifiers while using them. Keeping this in mind, we will summarise the findings from the experimental and phenomenological front, done so far, and we will provide the usage recommendations of these event classifiers. Eventually, to bring all collision systems on equal footing, we will also discuss the extension of these event topology classifiers to heavy-ion collisions, where the particle density in phase space is higher than $pp$ collisions. In this direction, the usefulness of machine learning is highlighted. Given that our focus in this review is to summarise event topology classifiers, rather than studying all possible theoretical/phenomenological models. We pick up two of the widely used event generators, i.e., PYTHIA 8 for $pp$ collisions and A Multi-Phase Transport (AMPT) [83, 84] for heavy-ion collisions, for some of the studies done here for a better understanding of event shapes and multihadron production dynamics. Wherever available, we will discuss the experimental results based on event classifiers.
@@
The paper is organized as follows. We begin with a brief introduction and motivation about the event shape observables in Sec. I. In Sec. II, we define each event classifier. Section III describes event generation methodology and event/track selection criteria used in this study according to experiments. Section IV discusses the results, especially highlighting selection biases, correlation among the classifiers, and particle ratios obtained in $pp$ collisions. Further, we extend the discussion of event shape studies to heavy-ion collisions and the application of machine learning techniques in Sec. V. The important findings of each event classifier are summarized in Sec. VI. Finally, Sec. VII presents the summary and provides a brief outlook.
II Event Classifiers

### v2

@@ Abstract
Event classifiers are the most fundamental observables to probe the event topology of hadronic and nuclear collisions at relativistic energies. Over the last five decades, significant progress has been made to establish suitable event classifiers to probe different physics processes occurring in elementary $e^{+}e^{-}$ to heavy-ion collisions in a broad range of center of mass energies. One of the major motivations to revisit event classifiers at the Large Hadron Collider (LHC) originates from the recent measurements of high multiplicity proton-proton collisions, which have revealed that these small collision systems exhibit features similar to the formation of quark-gluon plasma (QGP), traditionally believed to be only achievable in heavy nucleus-nucleus collisions at ultra-relativistic energies. To pinpoint the origin of these QGP-like phenomena with substantially reduced autocorrelation and selection biases, and to bring all collision systems on equal footing, along with charged-particle multiplicity, lately several event topology classifiers such as transverse sphericity, transverse spherocity, relative transverse activity classifier, and charged-particle flattenicity have been used extensively in experiments as well as in the phenomenological front. In addition, the infrared and collinear safety of event-shape observables makes them ideal for precision studies of jets and heavy-flavors at the LHC. In this review article, we summarise the motivation, scope, and practical use of these event-shape observables. The discussion integrates results and insights from all major LHC experiments, setting the stage for precision investigations for Run 3, Run 4, and future high luminosity upgrades of the LHC. In most cases, the event shape observables are found to be better probes in understanding the heavy-ion-like behavior seen at the LHC, while making a multi-differential study of multihadron production dynamics in hadronic and nuclear collisions.
keywords:
Event shape, ultra-relativistic collisions, Large Hadron Collider, quark-gluon plasma
1 Introduction
@@
To understand the infancy of our Universe and its evolution, the little bangs of heavy-ions are made in collider experiments like the Large Hadron Collider (LHC) at CERN, Switzerland, and the Relativistic Heavy-Ion Collider (RHIC) at BNL, USA, which reproduces the conditions that happened during the fraction of a second after the Big Bang. The collision of two nuclei produces a hot and dense system of deconfined quarks and gluons for a short time (order of a few fm/c or 10-23 sec). This deconfined and thermalized state of quarks and gluons is often referred to as quark-gluon plasma (QGP). Due to a very short timescale, it is impossible to probe QGP directly. However, several indirect signatures are proposed, such as the presence of strangeness enhancement, quarkonia suppression, jet quenching, collective flow, direct photon and dilepton production, which validate the presence of QGP. In proton-proton ($pp$) collisions, where the system size is significantly smaller than Pb-Pb, the main hard parton-parton scatterings produce high transverse momentum ($p_{\rm T}$) particles which can be described by the perturbative quantum chromodynamics (pQCD). In the hard regime of QCD, jets are produced from parton scatterings at large $p_{\rm T}$. As in this regime, the coupling constant ($\alpha_{s}$) is significantly small, precise pQCD calculations can be performed. The semi-hard and soft parton scatterings, which are not a part of the main hard scatterings, form the underlying event (UE) Diehl:2011yj , Field:2012kd . As the soft-QCD calculations are cumbersome due to a larger value of $\alpha_{s}$, the description of the UE is modeled using different QCD-inspired phenomenological approaches Christiansen:2024bhe . At the LHC energies, due to high parton densities and the composite nature of hadrons, more than one binary interaction among partons is possible in a single $pp$ collision, which is often referred to as multi-parton interactions (MPI) Sjostrand:1987su . The MPI phenomenon has been widely supported by experimental data ZEUS:2007xwd , D0:2009apj and is one of the key ingredients in the Monte Carlo (MC) event generators. Fig. 1 shows the pictorial representation of MPI in pp collisions. In such MPI-based models, the final state in $pp$ collisions is sensitive to the modeling of MPI. They are also sensitive to non-perturbative effects in the final state, such as color reconnection (CR), where the strings connecting the partonic endpoints from each individual scattering can color reconnect (Fig. 2) and color Ropes, where overlapping strings can act coherently to form a color rope having a large effective string tension Sjostrand:2014zea .
@@
Figure 2: Pictorial representation of Color Reconnection (CR) in the string fragmentation model: (a) two independent hard scatterings, (b) color reconnected string.
@@
Recent measurements of high multiplicity $pp$ collisions at the LHC energies have revealed that these systems exhibit features similar to quark-gluon plasma, such as the presence of radial and elliptic flow CMS:2016fnw , Grosse-Oetringhaus:2024bwr , ALICE:2019zfl , CMS:2010ifv , ATLAS:2015hzw , CMS:2015fgy , ATLAS:2016yzd , Busza:2018rrf , ALICE:2025bwp and enhanced production of strange hadrons with respect to that of charged pions ALICE:2017jyt , ALICE:2018pal , ALICE:2019avo , traditionally believed to be only achievable in ultra-relativistic nucleus-nucleus collisions. These measurements may point towards a common underlying physics mechanism across collision systems. However, no jet quenching signatures have been reported in small collision systems Nagle:2018nvi , ALICE:2023plt . The event generators currently available with relevant physics models fail to simultaneously reproduce all the observed phenomena in small collisions. This makes the origin of QGP-like phenomena observed in small systems like $pp$ and p-Pb unclear. Out of several theoretical models, the explanations from the MPI-based models with final state non-pQCD phenomena like CR Sjostrand:2014zea , rope hadronisation Bierlich:2015rha , Bierlich:2014xba , and string shoving Bierlich:2016vgw , inherited in PYTHIA8, are able to explain many of the QGP-like signatures in $pp$ collisions. MPI is a phenomenological approach and can not be measured in experiments. Thus, many experimental measurements are performed as a function of charged particle multiplicity in the final state, which has a strong correlation with MPI in MC models. On the other hand, the measurements with charged particle multiplicity show a stronger than linear increase of high-$p_{\rm T}$ particle yields in high-multiplicity (HM) $pp$ collisions relative to that of minimum bias (MB) events ALICE:2019dfi . This signifies a selection bias towards local fluctuations of choosing only hard $pp$ collisions. Such biases can be removed when the event selection is performed with multiplicity measured in a different pseudorapidity interval with respect to the observable of interest. However, it is found that such measurements are still biased by the hard processes contributing to the high-$p_{\rm T}$ particles ALICE:2022qxg . Thus, in small collision systems, such selection biases in the measurements hinder the search for the origin of QGP-like behavior.
@@
To pin down the origin of QGP-like phenomena with substantially diminished selection bias and to bring all the collision systems on an equal footing, along with charged-particle multiplicity ($N_{\rm{ch}}$), lately event shape observables such as transverse spherocity ($S_{0}$) and transverse sphericity ($S_{\rm T}$) have been inherited from the usage in $e^{+}e^{-}$ collisions since the late 1970s MovillaFernandez:1997fr , Donoghue:1979vi , Dasgupta:2003iq , Banfi:2004nk . Additionally, there are other proposed event shape observables which are used in heavy-ion collisions and jet studies such as third-jet resolution parameter, $y_{23}$ Bethke:2009ehn , OPAL:1989fep , JADE:1982xwg ; aplanarity, $A$; transverse thrust, $\tau_{T}$; its minor component, $T_{\perp,\rm min}$  Brandt:1964sa , ATLAS:2012tch , CMS:2011usu ; $\mathcal{F}$-parameter Banfi:2010xy , reduced flow vector, $q_{n}$ STAR:2002hbo , Fox-Wolfram moments Fox:1978vu , and event isotropy Cesarotti:2020hwb , Cesarotti:2020ngq , ATLAS:2023mny . Moreover, several new event shape observables have been recently constructed, for example, relative transverse activity classifiers such as $R_{\rm T}$, $R_{\rm T}^{\rm min}$, and $R_{\rm T}^{\rm max}$ developed by P. Skands et. al. and charged-particle flattenicity ($\rho_{\rm ch}$) developed by A. Ortiz et. al. to reduce the sensitivity to hard processes compared to the classifiers measured based only on charged-particle multiplicity Martin:2016igp , Bencedi:2021tst , Ortiz:2017jaz , Weber:2018ddv , Ortiz:2024ndh . In both experiments as well as in the phenomenological fronts, extensive explorations using the above-mentioned event-shape observables have been performed ALICE:2019dfi , ALICE:2022qxg , MovillaFernandez:1997fr , Donoghue:1979vi , Dasgupta:2003iq , Banfi:2004nk , Bethke:2009ehn , OPAL:1989fep , JADE:1982xwg , Brandt:1964sa , ATLAS:2012tch , Banfi:2010xy , STAR:2002hbo , Cesarotti:2020hwb , Cesarotti:2020ngq , ATLAS:2023mny , Martin:2016igp , Bencedi:2021tst , Ortiz:2017jaz , Ortiz:2024ndh , Weber:2018ddv , ALICE:2012cor , ATLAS:2012uka , ALICE:2011ac , ALICE:2019bdw , ALICE:2019mmy , ALICE:2022fnb , ALICE:2023yuk , ALICE:2023csm , ALICE:2023bga , ALICE:2023dei , ALICE:2024vaf , CMS:2025sws , CMS:2011usu . Also, we have explored the particle production dynamics as a function of transverse spherocity for light, strange Rath:2019izg , Tripathy:2019blo , Khuntia:2018qox , R:2023lku , Prasad:2024gqq and heavy flavor Deb:2020ige , Khatun:2019dml sectors in proton-proton collisions. The possibility of thermalization in small collision systems has been explored using spherocity as an event-shape observable in Ref. Deb:2019yjo , Deb:2020ezw . The global observables, identified particle production, and anisotropic flow coefficients are studied as a function of transverse spherocity in Pb–Pb collisions in Refs. Mallick:2020dzv , Tripathy:2025npe , Prasad:2021bdq , Mallick:2020ium , Mallick:2021hcs , Prasad:2022zbr , Prasad:2025ezg .
@@
In this review article, we extensively discuss the purpose, coverage, and usage guidelines of these event shape classifiers, which can be crucial in view of the upcoming measurements in the precision era at the LHC. Since 1970, there have been several event classifiers proposed and studied. Over the past decade, event topology classifiers have gradually become part of the common variables of collider physics. At the ATLAS and CMS collaborations, variables such as transverse sphericity, thrust, and heavy-jet mass are routinely used to separate hard and soft components of particle production and to characterize the underlying event. In the ALICE collaboration, similar techniques have opened new perspectives on collectivity, strangeness production, and multi-partonic interactions in small and large systems. Although each experiment has its own detector geometry and acceptance, the underlying ideas remain the same, i.e., capturing the shape and complexity of event dynamics in a way that is independent of model assumptions and free from biases. Their growing use across experiments shows how these observables now bridge the traditional boundaries between high-energy and heavy-ion physics, providing a unified framework for studying QCD matter at the LHC. In this article, we will mostly highlight the event classifiers which are popular in the early 21st century since the commencement of the LHC in view of the studies of small collision systems at the LHC. We will briefly discuss the most fundamental event classifier for the high-energy collisions, i.e., the charged-particle multiplicity measured both in mid and forward rapidities. Then the discussion will be followed by transverse sphero(i)city, which uses the information of charged-particle transverse momenta and the azimuthal angle for each particle. Then, we will discuss the event shape classifiers such as $R_{\rm T}$, $R_{\rm T}^{\rm min}$, and $R_{\rm T}^{\rm max}$, which take advantage of the underlying event and use the information of charged-particle multiplicity in the transverse region and azimuthal angle of the particles. Last but not least, the study will be focused on a relatively new event shape classifier, charged-particle flattenicity, which takes the information of final state particle multiplicity in different pseudorapidity and azimuthal angle ranges. Along with their novelty, the above-discussed event classifiers show significant correlation with $N_{\rm mpi}$, which will be extensively discussed in the next section. In most cases, the event shape observables are found to be better probes in understanding the heavy-ion-like behavior seen at the LHC. However, the coverage of event shape classifiers usually overlaps with each other. Thus, one needs to have a better understanding of these classifiers while using them. Keeping this in mind, we will summarise the findings from the experimental and phenomenological front, done so far, and we will provide the usage recommendations of these event classifiers. Additionally, to gain insight into the salient aspects of the pQCD domain, we discuss the event shape dependence of several hard probe observables associated with jets and heavy flavour production. By systematically investigating the event shape dependence of observables in both the soft and hard QCD sectors, we achieve a unified and comprehensive understanding of event shape dynamics in QCD. In this direction, we provide particular emphasis on jets and heavy flavour production, guided primarily by experimental data. Eventually, to bring all collision systems on equal footing, we will also discuss the extension of these event topology classifiers to heavy-ion collisions, where the particle density in phase space is higher than in $pp$ collisions. In this direction, the usefulness of machine learning is highlighted.
Wherever available, this review discusses experimental results based on event classifiers across LHC experiments, complemented by relevant observations from the pre-LHC era. Given that the primary aim of this review is to summarise the use of event topology classifiers rather than to exhaustively study all theoretical or phenomenological models, representative simulations from two widely used event generators are included to illustrate the physical interpretation of these observables: PYTHIA 8 for $pp$ collisions and A Multi-Phase Transport (AMPT) Lin:2004en , He:2017tla for heavy-ion collisions. These examples are intended to provide a qualitative understanding of how event-shape observables relate to multihadron production dynamics, rather than to present new model studies. PYTHIA is a pQCD-inspired general-purpose Monte Carlo event generator used to simulate relativistic hadronic, leptonic, and heavy-ion collisions from RHIC to LHC energies. In PYTHIA8, models and theories are incorporated for a number of physics aspects, such as parton distributions, hard and soft interactions, initial- and final-state parton showers, multiparton interactions, fragmentation, partonic and hadronic decays Corke:2010yf , Andersson:1983ia . A detailed explanation of all physics processes involved in PYTHIA8 can be found in Ref. Sjostrand:2007gs , manual . The two key features of PYTHIA8, i.e., color reconnection and rope hadronisation, are explicitly discussed in the Appendix 9.1 of the review. AMPT model, by contrast, is a hybrid model that combines both the initial partonic collisions and the final hadronic interactions effectively, as well as accounts for the transition between these two distinct phases of matter. It has four major components, namely, initial conditions, partonic interactions, conversion of partonic to hadronic matter through the hadronization mechanism, and hadron interactions. Appendix 9.3, highlights several additional MC models that are relevant to this review.
Moreover, in this review, the phenomena referred to as “QGP-like” in small collision systems are not treated as evidence for the formation of a distinct deconfined phase of partons. These observed phenomena emerge as a consequence of underlying multi-hadron production dynamics, which become prominent in high multiplicity environments. On the other hand, the event shape classifiers are used to select the sample with a specific final state topology, which is characterised by either high particle density in the final state or back-to-back jet topology. In addition, there are different physical processes that have distinct momentum space correlations, which can be captured by different event classifiers. These specific conditions, by the selection of specific event classifiers, can amplify the “QGP-like” signals irrespective of whether the origin is hydrodynamic, partonic, or purely hadronic in nature. Therefore, the observation of “QGP-like” signals in these selected events should be interpreted as the interplay among the manifestation of multihadron production dynamics, hadronization mechanisms, and global event geometry rather than referring only to the formation of QGP. Thus, in this review, the use of event shape classifiers is to classify and organize events based on the different multi-hadron production dynamics, while the question of ultimate physics interpretation of the observed mechanisms is open and model dependent.
@@
The paper is organized as follows. We begin with a brief introduction and motivation about the event shape observables in Sec. 1. In Sec. 2, we define each event classifier. Section 3 discusses the event shape classifiers from the pre-LHC era and provides a precursor for studies related to the LHC era. Section 4 discusses the soft probes with event classifiers, especially highlighting selection biases, correlation among the classifiers, and particle ratios obtained in $pp$ collisions. Section 5 briefly discusses the studies related to hard probes, i.e., jets and heavy-flavors with event shape classifiers. Further, we extend the discussion of event shape studies to heavy-ion collisions in Sec. 6. Section 7 provides a brief summary on the machine learning techniques to classify the events at LHC. Section 8 presents the summary and provides a brief outlook. The event generation methodology and event/track selection criteria used according to experiments are discussed in the appendix Sec. 9. To guide the readers, the summary of each section is mentioned in italic font at the end of the corresponding section.
2 Event Classifiers